OpenStudy (anonymous):

the perimeter of a rectangle equals the sum of the lengths of its four side. if the width of rectangle is 9 inches and its perimeter is 54 inches, find the length of the rectangle.

OpenStudy (jack1):



OpenStudy (jack1):

solve for L

OpenStudy (anonymous):

it be 54 inch

OpenStudy (jack1):

... im sorry? how?

OpenStudy (jack1):

P = sum of all side P = L + L + W + W P = 2L + 2W = 54 if W = 9 P = 2L + 2x9 = 54 2L + 18 = 54 are you ok with the above?

OpenStudy (jack1):

from here (2L + 18 = 54) you should be ok to solve for L, yeah?

OpenStudy (anonymous):

18

OpenStudy (jack1):

spot on dude